### v1

# Some lemmas on spectral radius of graphs: including an application

Abstract
For a graph $G$, the spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. In this paper, we give three lammas on $\rho(G)$ when $G$ contains a spanning complete bipartite graph. Moreover, an application was also included at the end.
Keywords: spectral radius; walk; cycle; planar graph.
2020 Mathematics Subject Classification: 05C50.
1 Introduction
All graphs considered in this paper are finite, undirected and simple. For a graph $G$, let $\overline{G}$ be its complement. The vertex set and edge set of $G$ are denoted by $V(G)$ and $E(G)$, respectively. For two vertices $u,v$, let $d_{G}(u,v)$ denote the distance between them (i.e., the minimum length of a path connecting them). We say $u\sim v$, if $u$ and $v$ are adjacent in $G$. For a certain integer $n$, let $K_{n},P_{n}$ and $C_{n}$ be the complete graph, the path and the cycle of order $n$, respectively. For any $\ell\geq 2$ graphs $G_{1},G_{2},...,G_{\ell}$, let $\cup_{1\leq i\leq\ell}G_{i}$ be the disjoint union of them. In particular, let $tG=\cup_{1\leq i\leq t}G$ for any integer $t\geq 1$. Let $G_{1}\vee G_{2}$ be the join of $G_{1}$ and $G_{2}$, obtained from $G_{1}\cup G_{2}$ by connecting all vertices of $G_{1}$ to all vertices of $G_{2}$. For any terminology used but not defined here, one may refer to [2].
Let $G$ be a graph with vertices $v_{1},v_{2},...,v_{n}$. The adjacency matrix of $G$ is an $n\times n$ matrix $(a_{ij})$, where $a_{ij}=1$ if $v_{i}\sim v_{j}$, and $a_{ij}=0$ otherwise. The spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. By the well-known Perron–Frobenius theorem, $\rho(G)$ has a non-negative eigenvector. A non-negative eigenvector corresponding to $\rho(G)$ is called a Perron vector of $G$. If $G$ is connected, any Perron vector of $G$ has positive entries. The spectral radius of a graph is closely related to its convergence property. In this paper, we give the following three lammas on spectral radius of graphs, which will be used.
Lemma 1.1
For integers $n_{1}\geq n_{2}+2\geq 1$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|\geq 0$. Then $\rho(G_{1})>\rho(G_{2})$.
Lemma 1.2
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq 1$ and $n_{1}\geq n_{2}+2$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 4}n_{s}\geq 100$. Then $\rho(G_{1})<\rho(G_{2})$.
Lemma 1.3
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq n_{5}\geq 1$ and $n_{1}\geq n_{2}+3$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 5}n_{s}\geq 220$. Then $\rho(G_{1})<\rho(G_{2})$.
The rest of the paper is organized as follows. In Section 2, we include several lemmas, which will be used later. The proofs of Lemmas 1.1, 1.2 and 1.3 are given in Section 3. In Section, we include an application on spectral radius of $C_{\ell}$-free planar graphs of order $n\geq 1.8\times 10^{17}$, where $5\leq\ell\leq n$.
2 Preliminaries
Let $G$ be a graph. For an integer $\ell\geq 1$, a walk of length $\ell$ in $G$ is an ordered sequence of vertices $v_{0}v_{1}\cdots v_{\ell}$, such that $v_{i}\sim v_{i+1}$ for any $0\leq i\leq\ell-1$. The vertex $v_{0}$ is called the starting vertex of the walk. For any $u\in V(G)$ and $\ell\geq 1$, let $w^{\ell}_{G}(u)$ be the number of walks of length $\ell$ starting at $u$ in $G$. Let $W^{\ell}(G)=\sum_{v\in V(G)}w^{\ell}_{G}(u)$. The following spectral formula was given in [9].
Theorem 2.1
([9]) For an integer $r\geq 2$, let $K_{n_{1},n_{2},...,n_{r}}$ be the complete $r$-partite graph of order $n$ with parts $V_{1},V_{2},...,V_{r}$, where $n_{i}=|V_{i}|\geq 1$ for $1\leq i\leq r$ and $\sum_{1\leq i\leq r}n_{i}=n$. For each $1\leq i\leq r$, let $H_{i}$ be a graph with vertex set $V(H_{i})\subseteq V_{i}$. Let $G$ be the graph obtained from $K_{n_{1},n_{2},...,n_{r}}$ by embedding the edges of $H_{i}$ into $V_{i}$ for all $1\leq i\leq r$. Set $\rho=\rho(G)$. Then
$\sum_{1\leq s\leq r}\frac{1}{1+\frac{n_{s}}{\rho}+\sum^{\infty}_{i=1}\frac{W^{i}(H_{s})}{\rho^{i+1}}}=r-1.$
The existence of the infinite series in Theorem 2.1 was guaranteed by the following two results. The first one is taken from Theorem 8.1.1 of [2].
Lemma 2.2
([2]) If $H$ is a subgraph of a connected graph $G$, then $\rho(H)\leq\rho(G)$, with equality if and only if $H=G$.
Lemma 2.3
([9]) Let $G$ be a graph of order $n$. Then $\sum^{\infty}_{k=1}\frac{W^{k}(G)}{x^{k}}$ exists for any $x>\rho(G)$.
3 Proofs of Lemmas 1.1, 1.2 and 1.3
Let $G$ be a graph. For a vertex $u$ and a walk $Q$ of $G$, we say that $Q$ crosses $u$, if $u$ is one of the vertices in the sequence of $Q$. For any integer $\ell\geq 1$ and $u,v\in V(G)$, let $W^{\ell}_{u,v}(G)$ be the number of walks $Q$ of length $\ell$ in $G$, such that $Q$ crosses $u$ and $v$. Let $P_{0}$ denote the empty graph.
Lemma 3.1
For integers $n_{1}\geq n_{2}+2\geq 3$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. Then $W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$ for any $\ell\geq 1$.
Proof: Let $G_{1}=P_{n_{1}}\cup P_{n_{2}}$ and $G_{2}=P_{n_{1}-1}\cup P_{n_{2}+1}$. For any given $\ell\geq 1$, we shall prove
$W^{\ell}(G_{1})-W^{\ell}(G_{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}}).$
Recall that $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$ in $G_{1}$. In $G_{2}$, denote $P_{n_{2}+1}=v_{1}v_{2}\cdots v_{n_{2}}v_{n_{2}+1}$. Let $X$ denote the set of walks $Q$ of length $\ell$ in $G_{1}$, such that $Q$ crosses $u_{1}$. Let $Y$ denote the set of walks $Q$ of length $\ell$ in $G_{2}$, such that $Q$ crosses $v_{1}$. Clearly, $W^{\ell}(G_{1})=|X|+W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}})$ and $W^{\ell}(G_{2})=|Y|+W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}})$. This means that
$W^{\ell}(G_{1})-W^{\ell}(G_{2})=|X|-|Y|.$
Now we divide $X$ in two parts: let $X_{1}$ denote the set of walks $Q$ of length $\ell$ in $G_{1}$, such that $Q$ crosses $u_{1}$ and $u_{n_{2}+2}$; let $X_{2}$ denote the set of walks $Q$ of length $\ell$ in $G_{1}$, such that $Q$ crosses $u_{1}$ and does not cross $u_{n_{2}+2}$. Clearly, $X=X_{1}\cup X_{2}$. By definitions of the sets $X_{1},X_{2},Y$, we see $|X_{2}|=|Y|$ and $|X_{1}|=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$. Thus $|X|-|Y|=|X_{1}|=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$. Hence, $W^{\ell}(G_{1})-W^{\ell}(G_{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$. This completes the proof. $\Box$
Lemma 3.2
Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
Proof: Let $P_{\infty}$ be a path of ”infinite” order, such that $u,v$ are two vertices of distance $\ell$ in $P_{\infty}$. That is,
$P_{\infty}=\cdots u_{\ell}\cdots u_{2}u_{1}uz_{1}z_{2}\cdots z_{\ell-1}vv_{1}v_{2}\cdots.$
(See Figure 1, where the ”semicircle” denotes the distance $\ell$ between $u$ and $v$.) Clearly, $W^{s}_{u,v}(G)\leq W^{s}_{u,v}(P_{\infty})$ for any $s\geq 1$. For any vertex $y\in P_{\infty}$ and integer $a\geq 1$, let $w^{a}_{uv}(y)$ be the number of walks $Q$ of length $a$ in $P_{\infty}$, such that $Q$ starts at $y$ and crosses $u$ and $v$.
Figure 1: The infinite path $P_{\infty}$
Claim 1. $w^{a}_{uv}(z_{i})\leq w^{a}_{uv}(u_{i})+w^{a}_{uv}(v_{\ell-i})$ for any $1\leq i\leq\ell-1$ and $a\geq 1$.
Proof of Claim 1. Let $a\geq 1$ be given. Clearly, $w^{a}_{uv}(z_{i})=0$ for $1\leq a\leq\ell-1$. Now assume $a\geq\ell$. Let $1\leq i_{0}\leq\ell-1$ be fixed. We will show that
$w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}}).$
Let $X$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ first before $v$. Let $Y$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ first before $u$. Clearly, $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|$.
Now we show that $|X|\leq w^{a}_{uv}(u_{i_{0}})$ and $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$. We show the first one. Let $D$ be the set of walks $Q$ starting at $u_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u,v$. We can define an injective map $\phi$ between $X$ and $D$. For any $Q\in X$, since $Q$ crosses $u$ first, we can denote
$Q=z_{i_{0}}z_{j_{1}}\cdots z_{j_{s}}uf_{1}f_{2}\cdots,$
where $1\leq j_{1},...,j_{s}\leq\ell-1$ and $f_{1},f_{2},$... are vertices of $P_{\infty}$. Define
$\phi(Q)=u_{i_{0}}u_{j_{1}}\cdots u_{j_{s}}uf_{1}f_{2}\cdots.$
It is easy to see that $\phi$ is injective. Hence, $|X|\leq|D|=w^{a}_{uv}(u_{i_{0}})$. Similarly, we can show that $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$. Then $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}})$. This finishes the proof of Claim 1. $\Box$
By Claim 1, we see
$\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})\leq\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(v_{i})$
for any $a\geq 1$. Clearly, $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})=0$ for any $i\geq\ell+1$ and $a\leq 2\ell$. Thus, for any $a\leq 2\ell$,
$W^{a}_{u,v}(P_{\infty})$ $=w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})$
$\leq 2(w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})).$
By symmetry, $w^{a}_{uv}(u)=w^{a}_{uv}(v)$ and $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})$ for any $a\geq 1$ and $i\geq 1$. Thus, for any $a\leq 2\ell$,
$W^{a}_{u,v}(P_{\infty})\leq 4(w^{a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})).$
Since $P_{\infty}$ is a path, each walk starting at $u_{i}$ corresponds to a $\pm$”-series:
$u_{i}+--+-+\cdots.$
(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $a\geq 0$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”). Thus,
$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=2\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
Then
$\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}$ $\leq\sum_{0\leq a\leq\ell}\frac{4}{x^{\ell+a}}(w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i}))$
$\leq\frac{8}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
$=\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
$\leq\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
$\leq\frac{8}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
$=\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
Note that $W^{s}_{u,v}(G)=0$ for any $s\leq\ell-1$ as $d_{G}(u,v)=\ell$. Then
$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Noting $s-3\leq 2(s-\ell-2)$ for $s\geq 2\ell+1$, we have
$\sum_{s\geq 2\ell+1}\frac{W^{s}_{u,v}(G)}{x^{s}}$ $\leq\sum_{s\geq 2\ell+1}\frac{2^{s}n}{x^{s}}$
$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}\frac{2^{s-3}}{x^{s-\ell-2}}~~~(using~x\geq\sqrt{n})$
$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}(\frac{4}{x})^{s-\ell-2}~~~(using~s-3\leq 2(s-\ell-2))$
$\leq\frac{8}{x^{\ell}}\sum_{j\geq 1}(\frac{4}{x})^{j}~~~(using~\ell\geq 2)$
$=\frac{1}{x^{\ell}}\frac{32}{x-4}~~~(using~x\geq 5).$
Consequently,
$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
This completes the proof. $\Box$
Now we are ready to prove the three lemmas.
Proof of Lemma 1.1. Let $G^{1}=P_{n_{1}}\cup P_{n_{2}}\cup T$ and $G^{2}=P_{n_{1}-1}\cup P_{n_{2}+1}\cup T$. Then $G_{j}=H\vee G^{j}$ for $j=1,2$. Set $\rho(G_{j})=\rho_{j}$ for any $1\leq j\leq 2$. We will prove $\rho_{1}>\rho_{2}$. By Theorem 2.1,
$\frac{1}{1+\frac{|H|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{\rho^{i}_{j}}}+\frac{1}{1+\frac{|G^{j}|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{\rho^{i}_{j}}}=1,$
where $1\leq j\leq 2$. For $1\leq j\leq 2$, let
$f(G_{j},x)=\frac{1}{1+\frac{|H|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{x^{i}}}+\frac{1}{1+\frac{|G^{j}|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{x^{i}}}.$
(Clearly, the domain of $f(G_{j},x)$ is of the form $x\geq a_{j}$ for some positive number $a_{j}$.) Clearly, $f(G_{j},x)$ is strictly increasing with respective to $x$ in its domain. Thus $\rho_{j}$ is the largest root of $f(G_{j},x)=1$.
In $G^{1}$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. For any $\ell\geq 1$, by Lemma 3.1,
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}}).$
This means that $W^{\ell}(G^{1})-W^{\ell}(G^{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\geq 0$. Clearly, $W^{n_{2}+1}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})=2$. It follows that
$\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{2}^{i}}>\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{2}^{i}}.$
This implies that
$f(G_{1},\rho_{2})<f(G_{2},\rho_{2}).$
Note that $f(G_{2},\rho_{2})=1$. So, $f(G_{1},\rho_{2})<1$. Recall that $f(G_{1},x)$ is strictly increasing in its domain, and $f(G_{1},\rho_{1})=1$. We must have $\rho_{1}>\rho_{2}$. This completes the proof. $\Box$
Proof of Lemma 1.2. Let $G^{1}=P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T$ and $G^{2}=P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T$. Then $G_{j}=H\vee G^{j}$ for $j=1,2$. Set $\rho(G_{j})=\rho_{j}$ for any $1\leq j\leq 2$. We will prove $\rho_{1}<\rho_{2}$. Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}}$ as a subgraph, we have
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
By Theorem 2.1,
$\frac{1}{1+\frac{|H|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{\rho^{i}_{j}}}+\frac{1}{1+\frac{|G^{j}|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{\rho^{i}_{j}}}=1,$
where $1\leq j\leq 2$. For $1\leq j\leq 2$, let
$f(G_{j},x)=\frac{1}{1+\frac{|H|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{x^{i}}}+\frac{1}{1+\frac{|G^{j}|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{x^{i}}}.$
Similar to Lemma 1.1, $f(G_{j},x)$ is strictly increasing with respective to $x$ in its domain, and $f(G_{j},\rho_{j})=1$.
In $G^{1}$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. In $G^{2}$, denote $P_{n_{3}+1}=v_{1}v_{2}\cdots v_{n_{3}+1}$. For any $\ell\geq 1$, by Lemma 3.1,
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}}),$
and
$W^{\ell}(P_{n_{3}+1}\cup P_{n_{4}-1})-W^{\ell}(P_{n_{3}}\cup P_{n_{4}})=W^{\ell}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1}).$
This means that
$W^{\ell}(G^{1})-W^{\ell}(G^{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1}).$
Note that $W^{i}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\leq W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})$ for any $i\geq 1$ by the definition of $W^{i}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$, since $n_{2}\geq n_{4}$. Thus,
$W^{\ell}(G^{1})-W^{\ell}(G^{2})\leq W^{\ell}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1}).$
Then
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})=\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})-W^{i}(G^{2})}{\rho_{1}^{i}}\leq\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})-W^{i}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})}{\rho_{1}^{i}}.$
Clearly, $W^{n_{4}}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})=2$ and $W^{n_{4}+1}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})\geq 2$. It follows that
$\sum^{\infty}_{i=1}\frac{W^{i}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})}{\rho_{1}^{i}}\geq\frac{2}{\rho_{1}^{n_{4}}}+\frac{2}{\rho_{1}^{n_{4}+1}}=\frac{1}{\rho_{1}^{n_{4}+1}}(2\rho_{1}+2).$
By Lemma 3.2 (by letting $\ell=n_{4}+1$ there),
$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}<e^{0.52}<1.7$. Thus,
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{4\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
[…]
Now consider $g(x)=5.8+\frac{6.8}{x-1}+\frac{16}{x-4}-x$ for $x\geq 10$. Clearly, $g(x)$ is decreasing for $x\geq 10$. By a calculation, $g(10)=-\frac{7}{9}<0$. Thus $g(x)<0$ for any $x\geq 10$. Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
Proof of Lemma 1.3. Let $G^{1}=P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T$ and $G^{2}=P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T$. Then $G_{j}=H\vee G^{j}$ for $j=1,2$. Set $\rho(G_{j})=\rho_{j}$ for any $1\leq j\leq 2$. We will prove $\rho_{1}<\rho_{2}$. Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}}$ as a subgraph, we have
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
By Theorem 2.1,
$\frac{1}{1+\frac{|H|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{\rho^{i}_{j}}}+\frac{1}{1+\frac{|G^{j}|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{\rho^{i}_{j}}}=1,$
where $1\leq j\leq 2$. For $1\leq j\leq 2$, let
$f(G_{j},x)=\frac{1}{1+\frac{|H|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{x^{i}}}+\frac{1}{1+\frac{|G^{j}|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{x^{i}}}.$
Clearly, $f(G_{j},x)$ is strictly increasing with respective to $x$ in its domain, and $f(G_{j},\rho_{j})=1$.
In $G^{1}$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. In $G^{2}$, denote $P_{n_{4}+1}=v_{1}v_{2}\cdots v_{n_{4}+1}$. For any $\ell\geq 1$, by Lemma 3.1,
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}})=W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})$
$W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\leq W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
Similarly,
$W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1})=W^{\ell}(P_{n_{1}-1}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{3}+1})$
$W^{\ell}_{u_{1},u_{n_{3}+2}}(P_{n_{1}-1})\leq W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
It follows that
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1})\leq 2W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
Also by Lemma 3.1,
$W^{\ell}(P_{n_{4}+1}\cup P_{n_{5}-1})-W^{\ell}(P_{n_{4}}\cup P_{n_{5}})=W^{\ell}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1}).$
It follows that
$W^{\ell}(G^{1})-W^{\ell}(G^{2})\leq 2W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1}).$
Thus,
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})\leq\sum^{\infty}_{i=1}\frac{2W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})}{\rho_{1}^{i}}.$
Clearly, $W^{n_{5}}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})=2$ and $W^{n_{5}+1}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})\geq 2$. It follows that
$\sum^{\infty}_{i=1}\frac{W^{i}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})}{\rho_{1}^{i}}\geq\frac{2}{\rho_{1}^{n_{5}}}+\frac{2}{\rho_{1}^{n_{5}+1}}=\frac{1}{\rho_{1}^{n_{5}+1}}(2\rho_{1}+2).$
By Lemma 3.2 (by letting $\ell=n_{5}+1$ there),
$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{220},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}<e^{0.405}<1.5$. Thus,
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$
[…]
Now consider $g(x)=11+\frac{12}{x-1}+\frac{32}{x-4}-x$ for $x\geq\sqrt{220}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{220}$. By a calculation, $g(\sqrt{220})=-0.01\cdots$. Thus $g(x)<0$ for any $x\geq\sqrt{220}$. So, $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
4 Spectral radius of $C_{\ell}$-free planar graphs
A graph is called planar, if it can be drawn in the plane without crossed edges. Recently, Tait and Tobin [6] proved that $K_{2}\vee P_{n-2}$ is the only extremal graph with the maximum spectral radius among all planar graphs of large order. This (asymptotically) solves a conjecture proposed by Boots and Royle [1], and independently by Cao and Vince [3]. It is also interesting for the researchers to determine the extremal graphs with the maximum spectral radius among all $C_{\ell}$-free planar graphs of order $n$, where $3\leq\ell\leq n$. The case $\ell=4$ was settled by Nikiforov [5], and Zhai and Wang [8]. Very recently, Fang, Lin and Shi [4] solved the case when $\ell=3$ and $\ell\ll n$. Xu, Lin and Fang [7] solved the case when $\frac{\ell}{n}$ is close to 1. They [7] proposed it as a problem for the general case. Armed with Lemma 1.1 and Lemma 1.2, using the typical spectral method we can solve this problem. That is, we have the following Theorem 4.1.
For a finite family of graphs $\mathcal{G}$, let ${\rm SPEX}(\mathcal{G})$ be the set of extremal graphs with the maximum spectral radius in $\mathcal{G}$.
Theorem 4.1
For integers $n\geq 1.8\times 10^{17}$ and $5\leq\ell\leq n$, let $\mathcal{G}_{n,\ell}$ be the set of all $C_{\ell}$-free planar graphs of order $n$. Then we have the following conclusions.
$(i)$ For $5\leq\ell<\frac{2n+5}{3}$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{\lceil\frac{\ell-3}{2}\rceil}\cup(a+1)P_{\lfloor\frac{\ell-3}{2}\rfloor}\cup P_{b})\right\}$, where $n-\ell+1=a\lfloor\frac{\ell-3}{2}\rfloor+b$ with $0\leq b<\lfloor\frac{\ell-3}{2}\rfloor$.
$(ii)$ For $\frac{2n+5}{3}\leq\ell\leq n$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{2\ell-n-4}\cup 2P_{n-\ell+1})\right\}$.
Declaration of competing interest
There is no conflict of interest.
Data availability statement
No data was used for the research described in the article.

### v2

@@
 In this paper, we give three lammas on $\rho(G)$ when $G$ contains a spanning complete bipartite graph.
-Moreover, an application was also included at the end.
-Keywords: spectral radius; walk; cycle; planar graph.
-2020 Mathematics Subject Classification: 05C50.
+Using these lemmas and typical spectral method, we characterized the unique extremal graph with the maximum spectral radius among all planar graphs of large order $n$ without a cycle of length $\ell$, where $5\leq\ell\leq n$.
+Keywords: spectral radius; planar graph; cycle; walk.
+2020 Mathematics Subject Classification: 05C50; 05C35.
 1 Introduction
@@
 Let $G$ be a graph with vertices $v_{1},v_{2},...,v_{n}$.
-The adjacency matrix of $G$ is an $n\times n$ matrix $(a_{ij})$, where $a_{ij}=1$ if $v_{i}\sim v_{j}$, and $a_{ij}=0$ otherwise.
+The adjacency matrix $A(G)$ of $G$ is an $n\times n$ matrix $(a_{ij})$, where $a_{ij}=1$ if $v_{i}\sim v_{j}$, and $a_{ij}=0$ otherwise.
 The spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix.
@@
 The spectral radius of a graph is closely related to its convergence property.
-In this paper, we give the following three lammas on spectral radius of graphs, which will be used.
+In this paper, we first give the following three lammas on spectral radius of graphs.
+(Let $P_{0}$ denote the empty graph.)
 Lemma 1.1
-For integers $n_{1}\geq n_{2}+2\geq 1$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|\geq 0$.
-Then $\rho(G_{1})>\rho(G_{2})$.
+For integers $n_{1}\geq n_{2}+2\geq 3$, let
+$G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup T)$
+and
+$G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup T),$
+where $H,T$ are two graphs with $|H|\geq 1$ and $|T|\geq 0$. Then $\rho(G_{1})>\rho(G_{2})$.
 Lemma 1.2
-For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq 1$ and $n_{1}\geq n_{2}+2$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 4}n_{s}\geq 100$.
-Then $\rho(G_{1})<\rho(G_{2})$.
+For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq 1$ and $n_{1}\geq n_{2}+2$, let
+$G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T)$
+and
+$G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T),$
+where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 4}n_{s}\geq 130$. Then $\rho(G_{1})<\rho(G_{2})$.
 Lemma 1.3
-For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq n_{5}\geq 1$ and $n_{1}\geq n_{2}+3$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 5}n_{s}\geq 220$.
-Then $\rho(G_{1})<\rho(G_{2})$.
+For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq n_{5}\geq 1$ and $n_{1}\geq n_{2}+3$, let
+$G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T)$
+and
+$G_{2}=H\vee(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T),$
+where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 5}n_{s}\geq 310$. Then $\rho(G_{1})<\rho(G_{2})$.
+A graph is called planar, if it can be drawn in the plane without crossed edges.
+Tait and Tobin [6] proved that $K_{2}\vee P_{n-2}$ is the only extremal graph with the maximum spectral radius among all planar graphs of large order $n$.
+This (asymptotically) solves a conjecture proposed by Boots and Royle [1], and independently by Cao and Vince [3].
+Recently, the researchers are also interested in determining the extremal graphs with the maximum spectral radius among all $C_{\ell}$-free planar graphs of large order $n$, where $3\leq\ell\leq n$.
+The case $\ell=4$ can be settled from the results of Nikiforov [5], and Zhai and Wang [8].
+Very recently, Fang, Lin and Shi [4] solved the case when $\ell=3$ and $5\leq\ell\ll n$.
+Xu, Lin and Fang [7] solved the case when $\frac{\ell}{n}$ is close to 1.
+They [7] proposed the general case as a problem.
+Armed with Lemma 1.1 and Lemma 1.2, using typical spectral method we can solve this problem.
+That is, we have the following Theorem 1.4.
+For a finite family of graphs $\mathcal{G}$, let ${\rm SPEX}(\mathcal{G})$ be the set of all extremal graphs with the maximum spectral radius in $\mathcal{G}$.
+Theorem 1.4
+For integers $n\geq 1.8\times 10^{17}$ and $5\leq\ell\leq n$, let $\mathcal{G}_{n,\ell}$ be the set of all $C_{\ell}$-free planar graphs of order $n$. Then we have the following conclusions.
+$(i)$ For $5\leq\ell<\frac{2n+5}{3}$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{\lceil\frac{\ell-3}{2}\rceil}\cup(a+1)P_{\lfloor\frac{\ell-3}{2}\rfloor}\cup P_{b})\right\}$, where $n-\ell+1=a\lfloor\frac{\ell-3}{2}\rfloor+b$ with $0\leq b<\lfloor\frac{\ell-3}{2}\rfloor$.
+$(ii)$ For $\frac{2n+5}{3}\leq\ell\leq n$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{2\ell-n-4}\cup 2P_{n-\ell+1})\right\}$.
 The rest of the paper is organized as follows.
@@
 The proofs of Lemmas 1.1, 1.2 and 1.3 are given in Section 3.
-In Section, we include an application on spectral radius of $C_{\ell}$-free planar graphs of order $n\geq 1.8\times 10^{17}$, where $5\leq\ell\leq n$.
+In Section 4, we give the proof of Theorem 1.4.
 2 Preliminaries
@@
 For any integer $\ell\geq 1$ and $u,v\in V(G)$, let $W^{\ell}_{u,v}(G)$ be the number of walks $Q$ of length $\ell$ in $G$, such that $Q$ crosses $u$ and $v$.
-Let $P_{0}$ denote the empty graph.
+Recall that $P_{0}$ denotes the empty graph.
 Lemma 3.1
@@
 $\Box$
+The following fact can be proved by induction on $\ell$.
+Fact 1. For integers $n\geq 4\ell\geq 0$, we have
+$\binom{n}{0}+\binom{n}{1}+\cdots+\binom{n}{\ell}\leq\frac{5}{4}\frac{n^{\ell}}{\ell!}$
 Lemma 3.2
@@
 Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$.
-Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
+Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
 Proof: Let $P_{\infty}$ be a path of ”infinite” order, such that $u,v$ are two vertices of distance $\ell$ in $P_{\infty}$. That is,
@@
 $u_{i}+--+-+\cdots.$
-(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $a\geq 0$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”).
-Thus,
-$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
-Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
-$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=2\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
+(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $0\leq a\leq\ell$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”).
+Thus, using Fact 1 as $\ell+a\geq 4\lfloor\frac{a-i}{2}\rfloor$,
+$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{0}+\binom{\ell+a}{1}+\cdots+\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{5}{4}\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
+Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{5}{4}\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
+$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times\frac{5}{4}\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=\frac{5}{2}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
 Then
 $\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}$ $\leq\sum_{0\leq a\leq\ell}\frac{4}{x^{\ell+a}}(w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i}))$
-$\leq\frac{8}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
-$=\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
-$\leq\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
-$\leq\frac{8}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
-$=\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
+$\leq\frac{10}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
+$=\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
+$\leq\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
+$\leq\frac{10}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
+$=\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
 Note that $W^{s}_{u,v}(G)=0$ for any $s\leq\ell-1$ as $d_{G}(u,v)=\ell$. Then
-$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
+$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
 Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Noting $s-3\leq 2(s-\ell-2)$ for $s\geq 2\ell+1$, we have
@@
 Consequently,
-$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
+$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
 This completes the proof. $\Box$
@@
 Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}}$ as a subgraph, we have
-$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
+$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
 By Theorem 2.1,
@@
 By Lemma 3.2 (by letting $\ell=n_{4}+1$ there),
-$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
-Since $\rho_{1}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}<e^{0.52}<1.7$. Thus,
+$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
+Since $\rho_{1}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{4}+1)+1}{\rho_{1}^{2}}}\leq e^{0.51}<1.67$.
+Thus,
 $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
-$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{4\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
+$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{5\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
 […]
-Now consider $g(x)=5.8+\frac{6.8}{x-1}+\frac{16}{x-4}-x$ for $x\geq 10$.
-Clearly, $g(x)$ is decreasing for $x\geq 10$.
-By a calculation, $g(10)=-\frac{7}{9}<0$.
-Thus $g(x)<0$ for any $x\geq 10$.
+Now consider $g(x)=8.5+\frac{7.5}{x-1}+\frac{16}{x-4}-x$ for $x\geq\sqrt{130}$.
+Clearly, $g(x)$ is decreasing for $x\geq\sqrt{130}$.
+By a calculation, $g(\sqrt{130})=-0.01\cdots<0$.
+Thus $g(x)<0$ for any $x\geq\sqrt{130}$.
 Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$.
@@
 Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}}$ as a subgraph, we have
-$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
+$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
 By Theorem 2.1,
@@
 $W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\leq W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
-Similarly,
+Similarly (denoting $P_{n_{1}-1}=u_{1}u_{2}\cdots u_{n_{1}-1}$),
 $W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1})=W^{\ell}(P_{n_{1}-1}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{3}+1})$
@@
 By Lemma 3.2 (by letting $\ell=n_{5}+1$ there),
-$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
-Since $\rho_{1}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{220},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}<e^{0.405}<1.5$. Thus,
+$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
+Since $\rho_{1}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{310},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{5}+\frac{6}{5})+\frac{4}{5}}{\rho_{1}^{2}}}<e^{0.403}<1.5$.
+Thus,
 $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
-$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$
+$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$
 […]
-Now consider $g(x)=11+\frac{12}{x-1}+\frac{32}{x-4}-x$ for $x\geq\sqrt{220}$.
-Clearly, $g(x)$ is decreasing for $x\geq\sqrt{220}$.
-By a calculation, $g(\sqrt{220})=-0.01\cdots$.
-Thus $g(x)<0$ for any $x\geq\sqrt{220}$.
+Now consider $g(x)=14+\frac{15}{x-1}+\frac{32}{x-4}-x$ for $x\geq\sqrt{310}$.
+Clearly, $g(x)$ is decreasing for $x\geq\sqrt{310}$.
+By a calculation, $g(\sqrt{310})=-0.35\cdots$.
+Thus $g(x)<0$ for any $x\geq\sqrt{310}$.
 So, $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$.
@@
 $\Box$
-4 Spectral radius of $C_{\ell}$-free planar graphs
-A graph is called planar, if it can be drawn in the plane without crossed edges.
-Recently, Tait and Tobin [6] proved that $K_{2}\vee P_{n-2}$ is the only extremal graph with the maximum spectral radius among all planar graphs of large order.
-This (asymptotically) solves a conjecture proposed by Boots and Royle [1], and independently by Cao and Vince [3].
-It is also interesting for the researchers to determine the extremal graphs with the maximum spectral radius among all $C_{\ell}$-free planar graphs of order $n$, where $3\leq\ell\leq n$.
-The case $\ell=4$ was settled by Nikiforov [5], and Zhai and Wang [8].
-Very recently, Fang, Lin and Shi [4] solved the case when $\ell=3$ and $\ell\ll n$.
-Xu, Lin and Fang [7] solved the case when $\frac{\ell}{n}$ is close to 1.
-They [7] proposed it as a problem for the general case.
-Armed with Lemma 1.1 and Lemma 1.2, using the typical spectral method we can solve this problem.
-That is, we have the following Theorem 4.1.
-For a finite family of graphs $\mathcal{G}$, let ${\rm SPEX}(\mathcal{G})$ be the set of extremal graphs with the maximum spectral radius in $\mathcal{G}$.
-Theorem 4.1
-For integers $n\geq 1.8\times 10^{17}$ and $5\leq\ell\leq n$, let $\mathcal{G}_{n,\ell}$ be the set of all $C_{\ell}$-free planar graphs of order $n$. Then we have the following conclusions.
-$(i)$ For $5\leq\ell<\frac{2n+5}{3}$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{\lceil\frac{\ell-3}{2}\rceil}\cup(a+1)P_{\lfloor\frac{\ell-3}{2}\rfloor}\cup P_{b})\right\}$, where $n-\ell+1=a\lfloor\frac{\ell-3}{2}\rfloor+b$ with $0\leq b<\lfloor\frac{\ell-3}{2}\rfloor$.
-$(ii)$ For $\frac{2n+5}{3}\leq\ell\leq n$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{2\ell-n-4}\cup 2P_{n-\ell+1})\right\}$.
+4 Proof of theorem 1.4
+For a graph $G$, let $e(G)=|E(G)|$.
+If $G$ is a planar graph of order $n\geq 2$, it is well-known that $e(G)\leq 3n-6$.
+Furthermore, $e(G)\leq 2n-4$ if $G$ is also bipartite.
+For $u\in V(G)$, let $d_{G}(u)$ denote the degree of $u$ in $G$.
+For $S\subseteq V(G)$, let $d_{S}(u)$ denote the number of neighbors of $u$ in $S$.
+Let $K_{s,t}$ be the complete bipartite graph of color orders $s$ and $t$.
+Recall that $\mathcal{G}_{n,\ell}$ is the set of all $C_{\ell}$-free planar graphs of order $n$.
+The following Lemma 4.1 is a similar result as Lemma 9 of [7].
+We need some changes, since we require this result for any $5\leq\ell\leq n$.
+(We point out that the bound $n\geq 1.8\times 10^{17}$ can be decreased using the current method. However, we will not do this, as it is not the main purpose of this paper.)
+Lemma 4.1
+For integers $n\geq 1.8\times 10^{17}$ and $5\leq\ell\leq n$, any graph $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$ contains a spanning subgraph $K_{2}\vee\overline{K_{n-2}}$.
+Proof: Let $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$.
+Recall that $G$ is a planar graph.
+We can assume that $G$ is already embedded in the plane.
+Since the resulting graph obtained by adding one suitable edge between two components of $G$, is still a planar graph without $C_{\ell}$ but has larger spectral radius by Lemma 2.2, we see that $G$ is connected as $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$.
+Let $\rho=\rho(G)$.
+(Let $\mathbf{x}^{T}$ denote the transpose of a vector $\mathbf{x}$.) By the Perron-Frobenius theorem, there exists a positive eigenvector $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})^{\mathrm{T}}$ corresponding to $\rho$ with $\max_{u\in V(G)}x_{u}=1$.
+Let $u^{\prime}\in V(G)$ with $x_{u^{\prime}}=1$.
+Clearly, $K_{2,n-2}$ is planar and $C_{\ell}$-free as $\ell\geq 5$, which implies $K_{2,n-2}\in\mathcal{G}_{n,\ell}$.
+Then
+$\rho\geq\rho(K_{2,n-2})=\sqrt{2n-4}.$
+Let $M=\{u\in V(G)~|~x_{u}\geq\frac{1}{10^{4}}\}$. Similar to Claims 10 and 11 of [7], we can show the following claim.
+Claim 1. $|M|\leq\frac{n}{10^{4}}$, and $d_{G}(u)\geq(x_{u}-\frac{8}{10^{4}})n$ for any $u\in M$.
+Choose a $u^{\prime\prime}\neq u^{\prime}$, such that $x_{u^{\prime\prime}}=\max_{u\in V(G)-\{u^{\prime}\}}x_{u}$. Similar to Claim 12 of [7], we can show the following claim.
+Claim 2. $x_{u^{\prime\prime}}\geq\frac{997}{1000}$.
+Recall that $x_{u^{\prime}}=1$ and $x_{u^{\prime\prime}}\geq\frac{997}{1000}$ by Claim 2.
+By Claim 1, we have $d_{G}(u^{\prime})\geq\frac{999n}{1000}$ and $d_{G}(u^{\prime\prime})\geq\frac{996n}{1000}$.
+Let $R=N_{G}(u^{\prime})\cap N_{G}(u^{\prime\prime})$ and $S=V(G)-(\{u^{\prime},u^{\prime\prime}\}\cup R)$.
+Then $|R|\geq d_{G}(u^{\prime})+d_{G}(u^{\prime\prime})-n\geq\frac{995n}{1000}$, and $|S|\leq(n-d_{G}(u^{\prime}))+(n-d_{G}(u^{\prime\prime}))\leq\frac{5n}{1000}$.
+Thus,
+$\frac{|R|}{|S|}\geq\frac{995}{5}=199.$
+Similar to Claim 13 of [7], we can show the following claim.
+Claim 3. $x_{u}\leq\frac{3}{100}$ for any $u\in V(G)-\{u^{\prime},u^{\prime\prime}\}$.
+The following claim is also needed in our proof.
+Claim 4. $x_{u}\geq\frac{1}{\rho}$ for any $u\in V(G)-\{u^{\prime},u^{\prime\prime}\}$.
+Proof of Claim 4.
+Suppose $x_{u}<\frac{1}{\rho}$ for some $u\in V(G)-\{u^{\prime},u^{\prime\prime}\}$.
+Then $1>\rho x_{u}=\sum_{v\in N_{G}(u)}x_{v}$.
+Let $G_{1}$ be the planar graph obtained from $G$ by deleting all edges incident with $u$ and adding an edge between $u$ and $u^{\prime}$.
+Clearly, $G_{1}$ is also $C_{\ell}$-free.
+Thus, $G_{1}\in\mathcal{G}_{n,\ell}$.
+But
+$\rho(G_{1})-\rho(G)$ $\geq\frac{\mathbf{x}^{\mathrm{T}}(A(G_{1})-A(G))\mathbf{x}}{\mathbf{x}^{\mathrm{T}}\mathbf{x}}$
+$=\frac{2}{\mathbf{x}^{\mathrm{T}}\mathbf{x}}(x_{u^{\prime}}-\sum_{v\in N_{G}(u)}x_{v})>0,$
+implying that $\rho(G_{1})>\rho(G)$. But this contradicts that $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$. $\Box$
+Figure 2: The local structure of the planar graph $G$
+In the planar embedding of $G$, we can assume that $G[\{u^{\prime},u^{\prime\prime}\}\cup R]$ is embedded as in Figure 2, where $u_{1},u_{2},\ldots,u_{|R|}\in R$ are the vertices in clockwise order around $u^{\prime\prime}$.
+Claim 5. $S$ is empty.
+Proof of Claim 5.
+Suppose to the contrary that $S$ is non-empty.
+Let $|S|=s\geq 1$.
+Note that $G$ is $K_{3,3}$-free, since $G$ is planar.
+Thus, for each $\in S, uu$ is adjacent to at most one of $u^{\prime}$ and $u^{\prime\prime}$, and is adjacent to at most 2 vertices in $R$.
+Moreover, the maximum degree of $G[R]$ is at most 2.
+Since $G[S]$ is also planar, there exists a vertex $v_{1}\in S$ with $d_{S}(v_{1})\leq 5$.
+Let $S_{0}=S$ and $S_{1}=S_{0}-\{v_{1}\}$.
+Repeat this process, we obtain a sequence of sets $S_{0},S_{1},\ldots,S_{s-1}$ such that $d_{S_{i-1}}(v_{i})\leq 5$ and $S_{i}=S_{i-1}-\{v_{i}\}$ for each $1\leq i\leq s-1$.
+Let $v_{s}$ be the unique vertex in $S_{s-1}$.
+Using Claim 3, for any $1\leq i\leq s$, we have
+$\sum\limits_{\begin{subarray}{c}w\sim v_{i}\\ w\in\{u^{\prime},u^{\prime\prime}\}\cup R\cup S_{i-1}\end{subarray}}x_{w}\leq 1+\sum\limits_{\begin{subarray}{c}w\sim v_{i}\\ w\in R\end{subarray}}x_{w}+\sum\limits_{\begin{subarray}{c}w\sim v_{i}\\ w\in S_{i-1}\end{subarray}}x_{w}\leq\frac{121}{100}<x_{u^{\prime}}+x_{u^{\prime\prime}}-\frac{7}{10}.$
+Then, noting $x_{v_{i}}\geq\frac{1}{\rho}$ by Claim 4,
+$\sum\limits_{i=1}^{s}x_{v_{i}}\left((x_{u^{\prime}}+x_{u^{\prime\prime}})-\sum\limits_{\begin{subarray}{c}w\sim v_{i}\\ w\in\{u^{\prime},u^{\prime\prime}\}\cup R\cup S_{i-1}\end{subarray}}x_{w}\right)\geq\frac{7}{10}\sum\limits_{i=1}^{s}x_{v_{i}}\geq\frac{7}{10}\frac{s}{\rho}\geq\frac{7}{10\rho}.$
+Note that in Figure 2, there are |R|1|R|-1 closed regions: $u^{\prime}u_{i}u^{\prime\prime}u_{i+1}u^{\prime}$ for all $1\leq i\leq|R|-1$.
+Recall that $\frac{|R|}{s}\geq 199$.
+Thus, there are at least $4$ consecutive regions containing no vertices in $S$.
+Without loss of generality, assume that such regions are $u^{\prime}u_{i}u^{\prime\prime}u_{i+1}u^{\prime}$ for $1\leq i\leq 4$.
+It follows that $u_{2},u_{3},u_{4}$ are not adjacent to the vertices in $S$, since $G$ is planar.
+Since $G$ is $K_{3,3}$-free, $u_{i}$ has at most two neighbors in $R$ for $2\leq i\leq 4$.
+Then,
+$\rho x_{u_{i}}=\sum_{u\in N_{G}(u_{i})}x_{u}=x_{u^{\prime}}+x_{u^{\prime\prime}}+\sum_{u\in N_{R}(u_{i})}x_{u}\leq d_{G}(u_{i})\leq 4$
+for each $2\leq i\leq 4$. Consequently, $x_{u_{i}}\leq\frac{4}{\rho}$, and hence
+$x_{u_{2}}x_{u_{3}}+x_{u_{3}}x_{u_{4}}\leq\frac{32}{\rho^{2}}.$
+Let $G_{2}$ be the graph obtained from $G$ by first deleting the edges $u_{2}u_{3}$, $u_{3}u_{4}$ and all the edges incident to at least one vertex in $S$, and then adding the edges $v_{i}u^{\prime}$ and $v_{i}u^{\prime\prime}$ crossing the closed region $u^{\prime}u_{2}u^{\prime\prime}u_{3}u^{\prime}$ for each $v_{i}\in S$.
+Clearly, $G_{2}$ is planar.
+Since $u_{3}$ and the vertices in $S$ are only adjacent to $u^{\prime}$ and $u^{\prime\prime}$ in $G_{2}$, each $C_{k}$ with $k\geq 5$ of $G_{2}$ contains at most one vertex of $S\cup\left\{u_{3}\right\}$.
+It follows that $G_{2}$ is also $C_{\ell}$-free, and hence $G_{2}\in\mathcal{G}_{n,\ell}$.
+Recall that $x_{u_{2}}x_{u_{3}}+x_{u_{3}}x_{u_{4}}\leq\frac{32}{\rho^{2}}$, and
+$\sum\limits_{i=1}^{s}x_{v_{i}}\left((x_{u^{\prime}}+x_{u^{\prime\prime}})-\sum\limits_{\begin{subarray}{c}w\sim v_{i}\\ w\in\{u^{\prime},u^{\prime\prime}\}\cup R\cup S_{i-1}\end{subarray}}x_{w}\right)\geq\frac{7}{10\rho}.$
+Then
+$\rho(G_{2})-\rho(G)$
+$\geq\frac{\mathbf{x}^{\mathrm{T}}(A(G_{2})-A(G))\mathbf{x}}{\mathbf{x}^{\mathrm{T}}\mathbf{x}}$
+$\geq\frac{2}{\mathbf{x}^{\mathrm{T}}\mathbf{x}}\left(\sum\limits_{i=1}^{s}x_{v_{i}}\left((x_{u^{\prime}}+x_{u^{\prime\prime}})-\sum\limits_{\begin{subarray}{c}w\sim v_{i}\\ w\in\{u^{\prime},u^{\prime\prime}\}\cup R\cup S_{i-1}\end{subarray}}x_{w}\right)-x_{u_{2}}x_{u_{3}}-x_{u_{3}}x_{u_{4}}\right)$
+$\geq\frac{2}{\mathbf{x}^{\mathrm{T}}\mathbf{x}}(\frac{7}{10\rho}-\frac{32}{\rho^{2}})$
+$>0~~~(using~\rho\geq\sqrt{2n-4}).$
+Hence, $\rho(G_{2})>\rho(G)$, contradicting that $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$. Therefore, $S$ is empty. $\Box$
+Claim 6. $u^{\prime}u^{\prime\prime}\in E(G)$.
+Proof of Claim 6.
+Suppose to the contrary that $u^{\prime}u^{\prime\prime}\notin E(G)$.
+By Claim 5, we see $V(G)=\left\{u^{\prime},u^{\prime\prime}\right\}\cup R$.
+Recall that $G[R]$ has maximum degree at most 2, as $G$ is $K_{3,3}$-free.
+Note that $G$ is $C_{\ell}$-free.
+Thus, $G[R]$ must be a linear forest with at least 3 components (otherwise $G$ has a Hamilton cycle).
+Then there exists some integer $1\leq i_{0}\leq n-2$ such that $u_{i_{0}}u_{i_{0}+1}\notin E(G)$.
+This implies that $u^{\prime}u_{i_{0}}u^{\prime\prime}u_{i_{0}+1}u^{\prime}$ is a face in the embedding of $G$.
+Let $G_{3}$ be the graph obtained from $G$ by adding the edge $u^{\prime}u^{\prime\prime}$ crossing the face $u^{\prime}u_{i_{0}}u^{\prime\prime}u_{i_{0}+1}u^{\prime}$.
+Clearly, $G_{3}$ is planar and $C_{\ell}$-free.
+But $\rho(G_{3})>\rho(G)$ by Lemma 2.2, contradicting that $G\in{\rm SPEX}(\mathcal{G}_{n,\ell})$.
+Therefore, $u^{\prime}u^{\prime\prime}\in E(G)$.
+From the above discussion, we can see that $G$ contains $K_{2}\vee\overline{K_{n-2}}$ as a spanning subgraph. This completes the proof. $\Box$
+The following observation is easy to see.
+Observation ([7]) Let $H=\cup_{i=1}^{t}P_{n_{i}}$, where $t\geq 2$ and $n_{1}\geq n_{2}\geq\cdots\geq n_{t}\geq 1$. Then $K_{2}\vee H$ is $C_{\ell}$-free if and only if $n_{1}+n_{2}\leq\ell-3$.
+Now we are ready to prove Theorem 1.4.
+Proof of Theorem 1.4.
+Let $G$ be a graph in ${\rm SPEX}(\mathcal{G}_{n,\ell})$.
+We will prove that $G$ is unique as stated in the theorem.
+By Lemma 4.1, $G$ contains $K_{2}\vee\overline{K_{n-2}}$ as a spanning subgraph.
+Let $u^{\prime},u^{\prime\prime}$ be the two dominating vertices of $G$.
+Let $R=V(G)-\left\{u^{\prime},u^{\prime\prime}\right\}$.
+Since $G$ is $K_{3,3}$-free, we see that $G[R]$ has maximum degree at most 2.
+Since $G$ is $C_{\ell}$-free, $G[R]$ must be a linear forest with $t\geq 3$ components.
+That is, $G[R]=\cup_{i=1}^{t}P_{n_{i}}$, where $t\geq 3,n_{1}\geq n_{2}\geq\cdots\geq n_{t}\geq 1$ and $n_{1}+n_{2}\leq\ell-3$.
+Clearly, $G=K_{2}\vee(\cup_{1\leq i\leq t}P_{n_{i}})$ and
+$\sum_{1\leq i\leq t}n_{i}=n-2.$
+Now we show that $n_{1}+n_{2}=\ell-3$. Otherwise, $n_{1}+n_{2}\leq\ell-4$. Let
+$G^{*}=K_{2}\vee(P_{n_{1}+1}\cup P_{n_{2}}\cup P_{n_{3}-1}\cup(\cup_{4\leq i\leq t}P_{n_{i}})).$
+Clearly, $G^{*}$ is planar and $C_{\ell}$-free (by the above observation). But $\rho(G^{*})>\rho(G)$ by Lemma 1.1, a contradiction. Hence,
+$n_{1}+n_{2}=\ell-3.$
+Now we prove the theorem by two cases depending on the values of $\ell$.
+$(i)$ $5\leq\ell<\frac{2n+5}{3}$. Let $a\geq 0,b$ be the integers, such that $n-\ell+1=a\lfloor\frac{\ell-3}{2}\rfloor+b$ with $0\leq b<\lfloor\frac{\ell-3}{2}\rfloor$.
+We first show that $t\geq 4$. Otherwise, $t\leq 3$. Then
+$\sum_{1\leq i\leq t}n_{i}\leq n_{1}+n_{2}+\frac{1}{2}(n_{1}+n_{2})\leq\frac{3}{2}(\ell-3)<\frac{3}{2}\frac{2n-4}{3}=n-2,$
+a contradiction. Hence $t\geq 4$.
+Now we show that $n_{1}-n_{2}\leq 1$. Otherwise, $n_{1}\geq n_{2}+2$. Let
+$G^{*1}=K_{2}\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup(\cup_{5\leq i\leq t}P_{n_{i}})).$
+Clearly, $G^{*1}$ is planar and $C_{\ell}$-free (by the above observation).
+But $\rho(G^{*1})>\rho(G)$ by Lemma 1.2, a contradiction.
+Hence, $n_{1}-n_{2}\leq 1$.
+It follows that $n_{1}=\lceil\frac{\ell-3}{2}\rceil$ and $n_{2}=\lfloor\frac{\ell-3}{2}\rfloor$.
+Now we show that $n_{3}=n_{4}=\cdots=n_{t-1}=\lfloor\frac{\ell-3}{2}\rfloor$. Otherwise, $n_{t-1}<n_{2}=\lfloor\frac{\ell-3}{2}\rfloor$. Let
+$G^{*2}=K_{2}\vee((\cup_{1\leq i\leq t-2}P_{n_{i}})\cup P_{n_{t-1}+1}\cup P_{n_{t}-1}).$
+Clearly, $G^{*2}$ is planar and $C_{\ell}$-free (by the above observation).
+But $\rho(G^{*2})>\rho(G)$ by Lemma 1.1, a contradiction.
+Hence, $n_{3}=n_{4}=\cdots n_{t-1}=\lfloor\frac{\ell-3}{2}\rfloor$.
+It follows that $t=a+3$ and $n_{t}=b$, since $n-\ell+1=a\lfloor\frac{\ell-3}{2}\rfloor+b$.
+Consequently, $G=K_{2}\vee(P_{\lceil\frac{\ell-3}{2}\rceil}\cup(a+1)P_{\lfloor\frac{\ell-3}{2}\rfloor}\cup P_{b})$, as desired.
+$(ii)$ $\frac{2n+5}{3}\leq\ell\leq n$. In this case, we have
+$\ell-2+\lfloor\frac{\ell-3}{2}\rfloor\geq\frac{2n+5}{3}-2+\lfloor\frac{n-2}{3}\rfloor\geq\frac{2n+5}{3}-2+\frac{n-4}{3}>n-2.$
+We first show that $t=3$.
+Otherwise, $t\geq 4$.
+If $n_{1}-n_{2}\geq 2$, we will obtain a contradiction using a similar discussion as that in $(i)$.
+Thus, we must have $n_{1}-n_{2}\leq 1$.
+It follows that $n_{1}=\lceil\frac{\ell-3}{2}\rceil$ and $n_{2}=\lfloor\frac{\ell-3}{2}\rfloor$, since $n_{1}+n_{2}=\ell-3$.
+We claim that $n_{3}=\lfloor\frac{\ell-3}{2}\rfloor$.
+Otherwise, $n_{t-1}\leq n_{3}<n_{2}=\lfloor\frac{\ell-3}{2}\rfloor$.
+Let
+$G^{*3}=K_{2}\vee((\cup_{1\leq i\leq 2}P_{n_{i}})\cup P_{n_{t-1}+1}\cup P_{n_{t}-1}).$
+Clearly, $G^{*3}$ is planar and $C_{\ell}$-free (by the above observation). But $\rho(G^{*3})>\rho(G)$ by Lemma 1.1, a contradiction. Hence, $n_{3}=\lfloor\frac{\ell-3}{2}\rfloor$. But then
+$\sum_{1\leq i\leq t}n_{i}\geq n_{1}+n_{2}+n_{3}+1=\ell-2+\lfloor\frac{\ell-3}{2}\rfloor>n-2,$
+a contradiction. Thus, we must have $t=3$. Then $G=K_{2}\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}})$.
+Now we show that $n_{2}=n_{3}$. Otherwise, $n_{2}\geq n_{3}+1$. Let
+$G^{*4}=K_{2}\vee(P_{n_{1}+1}\cup P_{n_{2}-1}\cup P_{n_{3}}).$
+Clearly, $G^{*4}$ is planar and $C_{\ell}$-free (by the above observation).
+But $\rho(G^{*4})>\rho(G)$ by Lemma 1.1, a contradiction.
+Hence, $n_{2}=n_{3}$.
+It follows that $n_{1}=2\ell-n-4$ and $n_{2}=n_{3}=n-\ell+1$, since $n_{1}+n_{2}+n_{3}=n-2$ and $n_{1}+n_{2}=\ell-3$.
+Consequently, $G=K_{2}\vee(P_{2\ell-n-4}\cup 2P_{n-\ell+1})$, as desired.
+This completes the proof.
+$\Box$
+Note. In this version, we correct some calculations, and include the detailed proof of Theorem 1.4.
 Declaration of competing interest

### v3

@@
 In Section 4, we give the proof of Theorem 1.4.
+Note. In this third version, we correct some calculation errors in Lemma 3.3.
 2 Preliminaries
@@
 $\Box$
-The following fact can be proved by induction on $\ell$.
-Fact 1. For integers $n\geq 4\ell\geq 0$, we have
-$\binom{n}{0}+\binom{n}{1}+\cdots+\binom{n}{\ell}\leq\frac{5}{4}\frac{n^{\ell}}{\ell!}$
+The following result will be used.
 Lemma 3.2
-Let $G$ be a path of order $n\geq 3$.
-Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$.
-Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
+For integers $n\geq 6m\geq 0$, we have
+$\binom{n}{0}4^{m}+\binom{n}{1}4^{m-1}+\cdots+\binom{n}{m}4^{0}\leq\frac{5}{2}\frac{n^{m}}{m!}$
+Proof: To prove the lemma, it is equivalent to show
+$\binom{n}{0}+\binom{n}{1}\frac{1}{4}+\binom{n}{2}\frac{1}{4^{2}}+\cdots+\binom{n}{m}\frac{1}{4^{m}}\leq\frac{5}{2}\frac{n^{m}}{4^{m}m!}.$
+We will prove this inequality by induction on $m$.
+It is easy to check for $m=0,1$.
+When $m=2$, we need to show $\binom{n}{0}+\binom{n}{1}\frac{1}{4^{1}}+\binom{n}{2}\frac{1}{4^{2}}\leq\frac{5}{2}\frac{n^{2}}{4^{2}2}$, where $n\geq 12$.
+This is equivalent to show $\frac{3}{2}n^{2}-3n-16\geq 0$, which is clearly true as $n\geq 12$.
+When $m=3$, we need to show $\binom{n}{0}+\binom{n}{1}\frac{1}{4^{1}}+\binom{n}{2}\frac{1}{4^{2}}+\binom{n}{3}\frac{1}{4^{3}}\leq\frac{5}{2}\frac{n^{3}}{4^{3}6}$, where $n\geq 18$.
+This is equivalent to show $3n^{3}-18n^{2}-172n-768\geq 0$, which is clearly true as $n\geq 18$.
+Now we suppose the inequality holds for $m=k$, where $k\geq 3$. We will prove the inequality for $m=k+1$. Now let $m=k+1$, implying $n\geq 6(k+1)$. Since the inequality holds for $m=k$, we have
+$\binom{n}{0}+\binom{n}{1}\frac{1}{4}+\binom{n}{2}\frac{1}{4^{2}}+\cdots+\binom{n}{k}\frac{1}{4^{k}}\leq\frac{5}{2}\frac{n^{k}}{4^{k}k!}.$
+Then
+$\binom{n}{0}+\binom{n}{1}\frac{1}{4}+\binom{n}{2}\frac{1}{4^{2}}+\cdots+\binom{n}{k}\frac{1}{4^{k}}+\binom{n}{k+1}\frac{1}{4^{k+1}}\leq\binom{n}{k+1}\frac{1}{4^{k+1}}+\frac{5}{2}\frac{n^{k}}{4^{k}k!}.$
+To prove the inequality for $m=k+1$, it suffices to show
+$\binom{n}{k+1}\frac{1}{4^{k+1}}+\frac{5}{2}\frac{n^{k}}{4^{k}k!}\leq\frac{5}{2}\frac{n^{k+1}}{4^{k+1}(k+1)!}.$
+This is equivalent to show
+$(n-1)(n-2)\cdots(n-k)\leq\frac{5}{2}n^{k-1}(n-4k-4).$
+It suffices to show
+$(n-k+1)(n-k)\leq\frac{5}{2}n(n-4k-4).$
+This is equivalent to show
+$n(3n-16k-22)\geq 2k^{2}-2k,$
+which is clearly true as $n\geq 6k+6$ and $k\geq 3$. Thus, the inequality also holds for $m=k+1$. By induction on $m$, this completes the proof. $\Box$
+Lemma 3.3
+Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then
+$\sum_{a\geq 1}\frac{W^{a}_{u,v}(G)}{x^{a}}\leq\frac{1}{x^{\ell}}(\frac{30x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}+\frac{32x}{x-8})$
+for any $x\geq\max\left\{\sqrt{n},9\right\}$.
 Proof: Let $P_{\infty}$ be a path of ”infinite” order, such that $u,v$ are two vertices of distance $\ell$ in $P_{\infty}$. That is,
@@
 Figure 1: The infinite path $P_{\infty}$
-Claim 1. $w^{a}_{uv}(z_{i})\leq w^{a}_{uv}(u_{i})+w^{a}_{uv}(v_{\ell-i})$ for any $1\leq i\leq\ell-1$ and $a\geq 1$.
-Proof of Claim 1. Let $a\geq 1$ be given. Clearly, $w^{a}_{uv}(z_{i})=0$ for $1\leq a\leq\ell-1$. Now assume $a\geq\ell$. Let $1\leq i_{0}\leq\ell-1$ be fixed. We will show that
-$w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}}).$
-Let $X$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ first before $v$.
-Let $Y$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ first before $u$.
-Clearly, $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|$.
-Now we show that $|X|\leq w^{a}_{uv}(u_{i_{0}})$ and $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$.
-We show the first one.
-Let $D$ be the set of walks $Q$ starting at $u_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u,v$.
+Claim 1.
+Let $1\leq a\leq\frac{3\ell}{2}$ be an integer.
+Then $w^{a}_{uv}(z_{i})\leq w^{a}_{uv}(u_{i})$ and $w^{a}_{uv}(z_{\ell-i})\leq w^{a}_{uv}(v_{i})$ for any $1\leq i<\frac{\ell}{2}$.
+Moreover, $w^{a}_{uv}(z_{\frac{\ell}{2}})\leq w^{a}_{uv}(u_{\frac{\ell}{2}})+w^{a}_{uv}(v_{\frac{\ell}{2}})$ for even $\ell$.
+Proof of Claim 1. Clearly, $w^{a}_{uv}(z_{i})=0$ for $1\leq a\leq\ell-1$. Now assume $\ell\leq a\leq\frac{3\ell}{2}$. Let $1\leq i_{0}<\frac{\ell}{2}$ be fixed. We will show that
+$w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}}).$
+Let $X$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ and $v$.
+Clearly, $w^{a}_{uv}(z_{i_{0}})=|X|$.
+Note that $1\leq i_{0}<\frac{\ell}{2}$ and $a\leq\frac{3\ell}{2}$.
+Each $Q\in X$ crosses $u$ first.
+Now we show that $|X|\leq w^{a}_{uv}(u_{i_{0}})$.
+Let $D$ be the set of walks $Q$ starting at $u_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ and $v$.
+Clearly, $w^{a}_{uv}(u_{i_{0}})=|D|$.
 We can define an injective map $\phi$ between $X$ and $D$.
@@
 Hence, $|X|\leq|D|=w^{a}_{uv}(u_{i_{0}})$.
-Similarly, we can show that $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$.
-Then $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}})$.
+That is, $w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}})$.
+Similarly, we can show that $w^{a}_{uv}(z_{\ell-i_{0}})\leq w^{a}_{uv}(v_{i_{0}})$.
+Now assume that $\ell\geq 2$ is even.
+Let $Y_{u}$ be the set of walks $Q$ starting at $z_{\frac{\ell}{2}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ first before $v$.
+Let $Y_{v}$ be the set of walks $Q$ starting at $z_{\frac{\ell}{2}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ first before $u$.
+Clearly, $w^{a}_{uv}(z_{\frac{\ell}{2}})=|Y_{u}|+|Y_{v}|$.
+Similar to the above discussion, we can show that $|Y_{u}|\leq w^{a}_{uv}(u_{\frac{\ell}{2}})$ and $|Y_{v}|\leq w^{a}_{uv}(v_{\frac{\ell}{2}})$.
+Hence $w^{a}_{uv}(z_{\frac{\ell}{2}})\leq w^{a}_{uv}(u_{\frac{\ell}{2}})+w^{a}_{uv}(v_{\frac{\ell}{2}})$.
 This finishes the proof of Claim 1.
@@
 By Claim 1, we see
-$\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})\leq\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(v_{i})$
-for any $a\geq 1$. Clearly, $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})=0$ for any $i\geq\ell+1$ and $a\leq 2\ell$. Thus, for any $a\leq 2\ell$,
-$W^{a}_{u,v}(P_{\infty})$ $=w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})$
-$\leq 2(w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})).$
-By symmetry, $w^{a}_{uv}(u)=w^{a}_{uv}(v)$ and $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})$ for any $a\geq 1$ and $i\geq 1$. Thus, for any $a\leq 2\ell$,
-$W^{a}_{u,v}(P_{\infty})\leq 4(w^{a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})).$
-Since $P_{\infty}$ is a path, each walk starting at $u_{i}$ corresponds to a $\pm$”-series:
+$\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})\leq\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(v_{i})$
+for any $1\leq a\leq\frac{3}{2}\ell$. Clearly, $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})=0$ for any $i>\frac{\ell}{2}$ and $a\leq\frac{3}{2}\ell$. Thus, for any $a\leq\frac{3}{2}\ell$,
+$W^{a}_{u,v}(P_{\infty})$ $=w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(v_{i})$
+$\leq 2\left(w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(v_{i})\right).$
+By symmetry, $w^{a}_{uv}(u)=w^{a}_{uv}(v)$ and $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})$ for any $a\geq 1$ and $i\geq 1$. Thus, for any $a\leq\frac{3\ell}{2}$,
+$W^{a}_{u,v}(P_{\infty})\leq 4(w^{a}_{uv}(u)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})).$
+Claim 2. Let $a,i$ be given integers with $\ell\leq a\leq\frac{3\ell}{2}$ and $1\leq i\leq\frac{\ell}{2}$. Then
+$w^{a}_{uv}(u_{i})\leq\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}$
+for even $a-\ell-i$, and
+$w^{a}_{uv}(u_{i})\leq\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}$
+for odd $a-\ell-i$.
+Proof of Claim 2. Since $P_{\infty}$ is a path, each walk starting at $u_{i}$ corresponds to a $\pm$”-series:
 $u_{i}+--+-+\cdots.$
-(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $0\leq a\leq\ell$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”).
-Thus, using Fact 1 as $\ell+a\geq 4\lfloor\frac{a-i}{2}\rfloor$,
-$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{0}+\binom{\ell+a}{1}+\cdots+\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{5}{4}\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
-Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{5}{4}\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
-$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times\frac{5}{4}\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=\frac{5}{2}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
+(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) Recall that $\ell\leq a\leq\frac{3\ell}{2}$ and $1\leq i\leq\frac{\ell}{2}$.
+Note that the distance between $u_{i}$ and $v$ is $\ell+i$.
+For any $\ell+i\leq j\leq a$, let $X_{j}$ be the set of walks $Q$ starting at $u_{i}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ at step $j$ (with smallest value).
+That is, $Q=u_{i}g_{1}g_{2}\cdots g_{a}$, where $g_{j}=v$ and $g_{k}\neq v$ for any $1\leq k\leq j-1$.
+Clearly, $j\equiv\ell+i(\hbox{\rm mod }~2)$.
+Moreover,
+$w^{a}_{uv}(u_{i})=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a}|$
+for even $a-\ell-i$, and
+$w^{a}_{uv}(u_{i})=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a-1}|$
+for odd $a-\ell-i$.
+For any $\ell+i\leq j\leq a$ and $j\equiv\ell+i(\hbox{\rm mod }~2)$, in the corresponding $\pm$”-series to walks $Q\in X_{j}$, there are exactly $\lfloor\frac{j-\ell-i}{2}\rfloor$ symbols ”-” in the first $j$ symbols $\pm$”.
+Hence,
+$|X_{j}|\leq\binom{j}{\lfloor\frac{j-\ell-i}{2}\rfloor}2^{a-j}.$
+Since $j\leq a\leq\lfloor\frac{3\ell}{2}\rfloor$, we have
+$|X_{j}|\leq\binom{\lfloor\frac{3\ell}{2}\rfloor}{\lfloor\frac{j-\ell-i}{2}\rfloor}2^{a-j}.$
+Let $m=\lfloor\frac{a-\ell-i}{2}\rfloor$. Clearly, $m\leq\frac{a-\ell-i}{2}\leq\frac{\ell}{4}$ as $a\leq\frac{3\ell}{2}$. Then $6m\leq\frac{3\ell}{2}$, implying $6m\leq\lfloor\frac{3\ell}{2}\rfloor$.
+For even $a-\ell-i$, using Lemma 3.2 as $6m\leq\lfloor\frac{3\ell}{2}\rfloor$,
+$w^{a}_{uv}(u_{i})$ $=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a}|$
+$\leq\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}2^{a-\ell-i}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}2^{a-\ell-i-2}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{2}2^{a-\ell-i-4}+\cdots+\binom{\lfloor\frac{3\ell}{2}\rfloor}{\frac{a-\ell-i}{2}}2^{0}$
+$=\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}4^{m}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}4^{m-1}+\cdots+\binom{\lfloor\frac{3\ell}{2}\rfloor}{m}4^{0}$
+$\leq\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{m}}{m!}$
+$=\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}.$
+For odd $a-\ell-i$, we have $m=\frac{a-1-\ell-i}{2}$. Using Lemma 3.2 as $6m\leq\lfloor\frac{3\ell}{2}\rfloor$,
+$w^{a}_{uv}(u_{i})$ $=|X_{\ell+i}|+|X_{\ell+i+2}|+\cdots+|X_{a-1}|$
+$\leq\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}2^{a-\ell-i}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}2^{a-\ell-i-2}+\binom{\lfloor\frac{3\ell}{2}\rfloor}{2}2^{a-\ell-i-4}+\cdots+\binom{\lfloor\frac{3\ell}{2}\rfloor}{\lfloor\frac{a-\ell-i}{2}\rfloor}2^{1}$
+$=2\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}2^{a-1-\ell-i}+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}2^{a-1-\ell-i-2}+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{2}2^{a-1-\ell-i-4}+\cdots+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{\lfloor\frac{a-\ell-i}{2}\rfloor}2^{0}$
+$=2\binom{\lfloor\frac{3\ell}{2}\rfloor}{0}4^{m}+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{1}4^{m-1}+\cdots+2\binom{\lfloor\frac{3\ell}{2}\rfloor}{m}4^{0}$
+$\leq\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{m}}{m!}$
+$=\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell-i}{2}\rfloor}}{\lfloor\frac{a-\ell-i}{2}\rfloor!}.$
+This finishes the proof of Claim 2. $\Box$
+Similar to Claim 2, we can show $w^{a}_{uv}(u)\leq\frac{5}{2}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell}{2}\rfloor}}{\lfloor\frac{a-\ell}{2}\rfloor!}$ for even $a$-$\ell$, and $w^{a}_{uv}(u)\leq\frac{5(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell}{2}\rfloor}}{\lfloor\frac{a-\ell}{2}\rfloor!}$ for odd $a$-$\ell$.
+Clearly, $w^{a}_{uv}(u_{i})=0$ for $i>a$-$\ell$.
+By Claim 2, for any $\ell\leq a\leq\frac{3\ell}{2}$,
+$w^{a}_{uv}(u)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i})$
+$=w^{a}_{uv}(u)+\sum_{1\leq i\leq a-\ell}w^{a}_{uv}(u_{i})$
+$\leq(5+\frac{5}{2})\times(1+\lfloor\frac{3\ell}{2}\rfloor+\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{2}}{2!}+\cdots+\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{\lfloor\frac{a-\ell}{2}\rfloor}}{\lfloor\frac{a-\ell}{2}\rfloor!})$
+$=\frac{15}{2}\sum_{0\leq i\leq\frac{a-\ell}{2}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}.$
 Then
-$\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}$ $\leq\sum_{0\leq a\leq\ell}\frac{4}{x^{\ell+a}}(w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i}))$
-$\leq\frac{10}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
-$=\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
-$\leq\frac{10}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
-$\leq\frac{10}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
-$=\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
+$\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{W^{a}_{u,v}(P_{\infty})}{x^{a}}$ $\leq\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{4}{x^{a}}(w^{a}_{uv}(u)+\sum_{1\leq i\leq\frac{\ell}{2}}w^{a}_{uv}(u_{i}))$
+$\leq\frac{30}{x^{\ell}}\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{1}{x^{a-\ell}}\sum_{0\leq i\leq\frac{a-\ell}{2}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}$
+$=\frac{30}{x^{\ell}}\sum_{0\leq i\leq\frac{\ell}{4}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}\sum_{\ell+2i\leq a\leq\frac{3\ell}{2}}\frac{1}{x^{a-\ell}}$
+$\leq\frac{30}{x^{\ell}}\sum_{0\leq i\leq\frac{\ell}{4}}\frac{(\lfloor\frac{3\ell}{2}\rfloor)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
+$\leq\frac{30}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}})^{i}$
+$=\frac{30}{x^{\ell}}\frac{x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}.$
 Note that $W^{s}_{u,v}(G)=0$ for any $s\leq\ell-1$ as $d_{G}(u,v)=\ell$. Then
-$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{10}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
-Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Noting $s-3\leq 2(s-\ell-2)$ for $s\geq 2\ell+1$, we have
-$\sum_{s\geq 2\ell+1}\frac{W^{s}_{u,v}(G)}{x^{s}}$ $\leq\sum_{s\geq 2\ell+1}\frac{2^{s}n}{x^{s}}$
-$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}\frac{2^{s-3}}{x^{s-\ell-2}}~~~(using~x\geq\sqrt{n})$
-$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}(\frac{4}{x})^{s-\ell-2}~~~(using~s-3\leq 2(s-\ell-2))$
-$\leq\frac{8}{x^{\ell}}\sum_{j\geq 1}(\frac{4}{x})^{j}~~~(using~\ell\geq 2)$
-$=\frac{1}{x^{\ell}}\frac{32}{x-4}~~~(using~x\geq 5).$
+$\sum_{1\leq a\leq\frac{3\ell}{2}}\frac{W^{a}_{u,v}(G)}{x^{a}}\leq\sum_{\ell\leq a\leq\frac{3\ell}{2}}\frac{W^{a}_{u,v}(P_{\infty})}{x^{a}}\leq\frac{30}{x^{\ell}}\frac{x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}.$
+Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Note that $a-5\leq 3(a-\ell-2)$ and $a-\ell-2\geq 0$, if $a>\frac{3\ell}{2}$ and $\ell\geq 2$. Thus,
+$\sum_{a>\frac{3\ell}{2}}\frac{W^{a}_{u,v}(G)}{x^{a}}$ $\leq\sum_{a>\frac{3\ell}{2}}\frac{2^{a}n}{x^{a}}$
+$\leq\frac{32}{x^{\ell}}\sum_{a>\frac{3\ell}{2}}\frac{2^{a-5}}{x^{a-\ell-2}}~~~(using~x\geq\sqrt{n})$
+$\leq\frac{32}{x^{\ell}}\sum_{a>\frac{3\ell}{2}}(\frac{8}{x})^{a-\ell-2}~~~(using~a-5\leq 3(a-\ell-2))$
+$\leq\frac{32}{x^{\ell}}\sum_{j\geq 0}(\frac{8}{x})^{j}~~~(using~a-\ell-2\geq 0)$
+$=\frac{32}{x^{\ell}}\frac{x}{x-8}~~~(using~x\geq 9).$
 Consequently,
-$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{10x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
+$\sum_{a\geq 1}\frac{W^{a}_{u,v}(G)}{x^{a}}\leq\frac{1}{x^{\ell}}(\frac{30x}{x-1}e^{\frac{\lfloor\frac{3\ell}{2}\rfloor}{x^{2}}}+\frac{32x}{x-8}).$
 This completes the proof. $\Box$
@@
 Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}}$ as a subgraph, we have
-$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
+$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{\sqrt{2000},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
 By Theorem 2.1,
@@
 By Lemma 3.3 (by letting $\ell=n_{4}+1$ there),
-$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
-Since $\rho_{1}\geq\max\left\{\sqrt{130},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{4}+1)+1}{\rho_{1}^{2}}}\leq e^{0.51}<1.67$.
+$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{30\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32\rho_{1}}{\rho_{1}-8}).$
+Since $\rho_{1}\geq\max\left\{\sqrt{2000},\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{\sqrt{2000},\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{\frac{3}{2}(n_{4}+1)}{\rho_{1}^{2}}}\leq e^{\frac{\frac{3}{2}(n_{4}+1)}{4n_{4}+2}}\leq e^{0.5}<1.65$.
 Thus,
 $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
-$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{5\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
+$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{15\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16\rho_{1}}{\rho_{1}-8}-\rho_{1}-1)$
 […]
-Now consider $g(x)=8.5+\frac{7.5}{x-1}+\frac{16}{x-4}-x$ for $x\geq\sqrt{130}$.
-Clearly, $g(x)$ is decreasing for $x\geq\sqrt{130}$.
-By a calculation, $g(\sqrt{130})=-0.01\cdots<0$.
-Thus $g(x)<0$ for any $x\geq\sqrt{130}$.
+[…]
+Now consider $g(x)=40+\frac{25}{x-1}+\frac{128}{x-8}-x$ for $x\geq\sqrt{2000}$.
+Clearly, $g(x)$ is decreasing for $x\geq\sqrt{2000}$.
+By a calculation, $g(\sqrt{2000})=-0.66\cdots<0$.
+Thus $g(x)<0$ for any $x\geq\sqrt{2000}$.
 Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$.
@@
 Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}}$ as a subgraph, we have
-$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
+$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{6500},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
 By Theorem 2.1,
@@
 By Lemma 3.3 (by letting $\ell=n_{5}+1$ there),
-$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
-Since $\rho_{1}\geq\max\left\{\sqrt{310},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{310},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}=e^{\frac{(2n_{5}+\frac{6}{5})+\frac{4}{5}}{\rho_{1}^{2}}}<e^{0.403}<1.5$.
+$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{30\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32\rho_{1}}{\rho_{1}-8}).$
+Since $\rho_{1}\geq\max\left\{\sqrt{6500},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{6500},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{\frac{3}{2}(n_{5}+1)}{\rho_{1}^{2}}}\leq e^{\frac{\frac{3}{2}(n_{5}+1)}{5n_{5}+3}}<e^{0.4}<1.5$.
 Thus,
 $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
-$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{10\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$
+$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{30\rho_{1}}{\rho_{1}-1}e^{\frac{\frac{3}{2}(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32\rho_{1}}{\rho_{1}-8}-\rho_{1}-1)$
 […]
-Now consider $g(x)=14+\frac{15}{x-1}+\frac{32}{x-4}-x$ for $x\geq\sqrt{310}$.
-Clearly, $g(x)$ is decreasing for $x\geq\sqrt{310}$.
-By a calculation, $g(\sqrt{310})=-0.35\cdots$.
-Thus $g(x)<0$ for any $x\geq\sqrt{310}$.
+[…]
+Now consider $g(x)=76+\frac{45}{x-1}+\frac{256}{x-8}-x$ for $x\geq\sqrt{6500}$.
+Clearly, $g(x)$ is decreasing for $x\geq\sqrt{6500}$.
+By a calculation, $g(\sqrt{6500})=-0.53\cdots$.
+Thus $g(x)<0$ for any $x\geq\sqrt{6500}$.
 So, $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$.
@@
 Therefore, $u^{\prime}u^{\prime\prime}\in E(G)$.
+$\Box$
 From the above discussion, we can see that $G$ contains $K_{2}\vee\overline{K_{n-2}}$ as a spanning subgraph. This completes the proof. $\Box$
@@
 $\Box$
-Note. In this version, we correct some calculations, and include the detailed proof of Theorem 1.4.
 Declaration of competing interest